volume of gas at stp is 1.12 cc calculate no of molecules in it ?
 

Grade:12th pass

1 Answers

kanchan
6 years ago
1 l = 1000 cc and at stp no of particle is 22.4 l
22.4*10= 1 mole 
1.12*10-7 = 1.12*10-7/22.4*103 = 1/20 * 10 -10 = 5*10-9 mole 
and we know that 1 mole = 6.02*1023
3.01*1012